40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Feb 3 20:19:38 2012, rana, Update, SUS, Oplev QPD Sum Trends are suspicious a.png
    Reply  Mon Feb 6 11:07:21 2012, steve, Update, SUS, Oplev QPD Sum Trends are suspicious oplevsums.png
Message ID: 6256     Entry time: Mon Feb 6 11:07:21 2012     In reply to: 6253
Author: steve 
Type: Update 
Category: SUS 
Subject: Oplev QPD Sum Trends are suspicious 

Quote:

The attached trend shows a problem with the QPD sums.

Why are ETMX and ITMX so much lower than ETMY and ITMY? Are the laser's dying? Or is it the gain inside the QPD? Or the reflectivity of the coatings?

Steve - please check on Monday the laser powers and the ETM/ITM reflectivity for HeNe lasers. Maybe we have to increase the transimpedance gain in the heads.

ETMX and ETMY have 0.2 mW returning to their QPDs........so the gain must lower at  ETMX

ITMX laser 1103P has only 0.67 mW output and 0.025 mW returning to the QPD.

ITMX and ETMX oplev lanching paths  have lenses without AR coating. This is my fault. I will buy them.

          ETMX   0.2 mW           900 counts

         ITMX     0.025           1300

         SRM      0.04            2600

          BS      0.05            3500

          PRM     0.06            4000

         ETMY     0.2             9000

        ITMY      0.3            14500

Attachment 1: oplevsums.png  58 kB  | Hide | Hide all
oplevsums.png
ELOG V3.1.3-